OpenStudy (anonymous):

your school sells tickets for its winter concert. Student tickets are $5 and adult tickets are $10. If your school sells 85 tickets and makes $600, how many of each ticket did they sell?

OpenStudy (anonymous):

This problem involves simultaneous equations. To get you started on the right path, setup the model for the school concert.

OpenStudy (anonymous):

First, let x represent student tickets and y are the adult tickets. What do you know about x and y?

OpenStudy (anonymous):

I started with x+y=600, then ..600/15= 40... but I can't connect the things..

OpenStudy (anonymous):

When x and y are added, you should have the sum persons going to the event. 600 is the dollar amount collected. So re-think your initial equations.

OpenStudy (anonymous):

did the same.. thanks ....

OpenStudy (ness9630):

Here, you need to set up two equations: \[x+y=85\] \[5x+10y=600\] Where x=student tickets and y=adult tickets. Can you handle it from here?